OpenStudy (anonymous):

solve: 2(cos^2)x= 3(cos^2)x

OpenStudy (anonymous):

simply as that...

sam (.sam.):

\[2\cos^2(x)= 3\cos^2(x)\]?

OpenStudy (anonymous):

$$2\cos^2x=3\cos^2x\\3\cos^2x-2\cos^2x=0\\(\cos^2x)(3-2)=0\\\cos^2x=0\implies\cos x=0$$What angles do you know have a cosine of \(0\)? Which angles project no shadow on the x-axis?